You are considering opening a new plant. The plant will cost $ 101.8 million upfront and will take one year to build. After​ that, it is expected to produce profits of $ 30.5 million at the end of every year of production. The cash flows are expected to last forever. Calculate the NPV of this investment opportunity if your cost of capital is 7.9 % . Should you make the​ investment? Calculate the IRR. Does the IRR rule agree with the NPV​ rule?

NPV= 255.9 milliom

IRR=?
Answer #1

NPV = CF/r - Initial investment

NPV = 30.5/0.079 - 101.8

NPV = $284.2759493671 million

Yes, you should invest in this new plant because NPV is positive.

IRR is the rate at which NPV is zero.

NPV = CF/r - Initial investment

0 = 30.5/IRR - 101.8

30.5/IRR = 101.8

IRR = 30.5/101.8

IRR = 0.2996070727

IRR = 29.96070727%

IRR > cost of capital, so IRR is also suggesting investing in this new plant.

Yes, the IRR rule agrees with the NPV rule.

Can you please upvote? Thank You :-)